A force of 6N extends a spring of natural length 0.6m by 0.02m .What will be the length of the spring when the applied force is 30N.
1
Expert's answer
2020-08-02T15:07:00-0400

According to the Hooke's law, the force "F" and extension "\Delta x" are connected in a following manner:


"F = k\Delta x"

where "k" is a spring constant. Expressing "k" using the first measurement, get:


"k = \dfrac{F_1}{\Delta x_1} = \dfrac{6N}{0.02m} = 300 N/m"

Next, let's find the extension for the second force:


"\Delta x_2 = \dfrac{F_2}{k} = \dfrac{30N}{300 N/m} = 0.1m"

The total length is the sum of the natural length 0.6m and the extension:


"l_2 = l_0 + \Delta x_2 = 0.6m + 0.1m = 0.7m"

Answer. 0.7 m.